# Learning Complex Motion Plans using Neural ODEs with Safety and Stability Guarantees

### v3

@@ IV Learning Nominal Motion Plans using Neural ODEs
to learn the parameters $\theta$, where the predictions of the state $\hat{x}_{i}(t_{k})$ are obtained by integrating (3) with initial condition $\hat{x}_{i}(t_{1})=x_{i}(t_{1}),\ \forall\ i\in\{1,2,\ldots,M\}$. [^1] A binomial checkpoint method is used during training for solving (4) as implemented in [27]. In contrast to previous work [16, 20] which learns a map $\hat{f}(\cdot)$ using labeled data $\{x(t),\dot{x}(t)\}$, we do not assume access to velocity measurements as they are often not easily collected and/or noisy [28, 29]. Fitting a map to noisy measurements lead to aggressive trajectories at inference that are not desirable for the low-level controller. From our results presented in Fig. 1 and Section VI, we observe that the NODE model generates smooth trajectories utilizing only state variables $x(t)$ to learn $f_{\theta}$ and not their derivatives $\dot{x}(t)$. While such a NODE-based vector field will behave reliably near the training data, unanticipated disturbances or obstacles during deployment might deviate the robot to regions of the state-space where the learned vector field is unreliable, as shown in Fig. 0(a). We present a method that computes a correction term to ensure robust and safe tracking of the learned target trajectory.
@@ V-A Preliminaries on Safety and Stability
$K(x):=\{u\in\mathcal{U}:\nabla_{x}B(x)^{\top}\left(g(x)+h(x)u\right)\geq-\gamma(B(x))\}.\vspace{-5pt}$ (8)
@@ V-B2 Safety using Control Barrier Functions
$\nabla_{x}B(x)^{\top}\left(f_{\theta}(x)+u(x)\right)\geq-\gamma(B(x)),\ \forall\ x\in\mathcal{X}\vspace{-5pt}$ (13)
@@ VI-B Periodic trajectories
We validate our approach on handwritten 2D periodic motions of the letters I, R, O and S given in [17] and 3D periodic trajectories that encode three wiping tasks as given in Figs. 4(e)5(a), and 5(b). We compare our method with Imitation Flow (IFlow) [17] and a Gaussian Process (GP) [11] based approach, since SEDS and LPV-DS compared in Section VI-A model trajectories that converge only to a single attractor. IFlow is based on normalizing flows that learns stable motions using prior knowledge on whether the underlying dynamics of the demonstrations has a single attractor or a limit cycle, but our approach requires no prior knowledge. We present the DTW distance in Fig. 4(a), training time in Fig. 4(b) and execution time in Fig. 4(c). The execution time is the computation time for a single forward pass of the model to generate the entire trajectory from a given initial point and the time span of the demonstration. We also compare the trajectory reproductions for the R shape in Fig. 4(d), and for the spiral wiping task by the Franka robot arm in Fig. 4(e). The IFlow approach is not able to learn the complex motion for spiral wiping, but our NODE approach learns with high accuracy and lesser computation time. The execution time comparison in Fig. 4(c) is plotted in $\log$ scale and we note that our approach (NODE) has much lesser execution time, which is important for real-time robot experiments. Although the GP based method [11] learns complex trajectories with comparable accuracy and training time to NODE, the execution time is much smaller and they rely on time inputs for desired roll outs with no capability to generate safe and stable motion plans. All computations are performed on Google Colab.
@@ VI-C Robotic experiments
We validate our approach on the Franka robot arm performing complex periodic motions: wiping a mannequin with a towel (Fig. 5(a)) and wiping a white board with an eraser (Fig. 5(b)). We use the passive DS impedance controller [36]. We used $2$ demonstrations for the mannequin task, and $3$ demonstrations for the board wiping task. Each demonstration had between $300$ and $600$ data samples. The average training time (offline) is $3-6$ minutes for each task on Google Colab. The obstacle shown in Fig. 5(b) at $t=2$ has markers on it that are tracked in real-time by OptiTrack motion capture system. We observe that the robot tracks the desired nominal trajectories while remaining compliant to human interaction, robust to perturbations, and safe with respect to unforeseen dynamic obstacles. The online motion planning module runs at 1KHz that solves (14) and implements Algorithm 1. We include the stirring task, the effect of Algorithm 1 and some preliminary results for full pose (position in $\mathbb{R}^{3}$ and orientation in $\mathcal{SO}(3)$) motion planning in https://sites.google.com/view/lfd-neural-ode/home and the video submission.

### v4

@@ IV Learning Nominal Motion Plans using Neural ODEs
[…]
@@ V-A Preliminaries on Safety and Stability
$K(x):=\{u\in\mathcal{U}:\nabla_{x}B(x)^{\top}\left(g(x)+h(x)u\right)\geq-\gamma(B(x))\}.\vskip-5.0pt$ (8)
@@ V-B2 Safety using Control Barrier Functions
$\nabla_{x}B(x)^{\top}\left(f_{\theta}(x)+u(x)\right)\geq-\gamma(B(x)),\ \forall\ x\in\mathcal{X}\vskip-5.0pt$ (13)
@@ VI-B Periodic trajectories
We validate our approach on handwritten 2D periodic motions of the letters I, R, O and S given in [17] and 3D periodic trajectories that encode three wiping tasks as given in Figs. 5(e)6(a), and 6(b). We compare our method with Imitation Flow (IFlow) [17] and a Gaussian Process (GP) [11] based approach, since SEDS and LPV-DS compared in Section VI-A model trajectories that converge only to a single attractor. IFlow is based on normalizing flows that learns stable motions using prior knowledge on whether the underlying dynamics of the demonstrations has a single attractor or a limit cycle, but our approach requires no prior knowledge. We present the DTW distance in Fig. 5(a), training time in Fig. 5(b) and execution time in Fig. 5(c). The execution time is the computation time for a single forward pass of the model to generate the entire trajectory from a given initial point and the time span of the demonstration. We also compare the trajectory reproductions for the R shape in Fig. 5(d), and for the spiral wiping task by the Franka robot arm in Fig. 5(e). The IFlow approach is not able to learn the complex motion for spiral wiping, but our NODE approach learns with high accuracy and lesser computation time. The execution time comparison in Fig. 5(c) is plotted in $\log$ scale and we note that our approach (NODE) has much lesser execution time, which is important for real-time robot experiments. Although the GP based method [11] learns complex trajectories with comparable accuracy and training time to NODE, the execution time is much longer and they rely on time inputs for desired roll outs with no capability to generate safe and stable motion plans. All computations are performed on Google Colab.
@@ VI-C Robotic experiments
We validate our approach on the Franka robot arm performing complex periodic motions: wiping a mannequin with a towel (Fig. 6(a)) and wiping a white board with an eraser (Fig. 6(b)). We use the passive DS impedance controller [36]. We used $2$ demonstrations for the mannequin task, and $3$ demonstrations for the board wiping task. Each demonstration had between $300$ and $600$ data samples. The average training time (offline) is 363-6 minutes for each task on Google Colab. The obstacle shown in Fig. 6(b) at $t=2$ has markers on it that are tracked in real-time by OptiTrack motion capture system. We observe that the robot tracks the desired nominal trajectories while remaining compliant to human interaction, robust to perturbations, and safe with respect to unforeseen dynamic obstacles. The online motion planning module runs at 1KHz that solves (14) and implements Algorithm 1. We include the stirring task, the effect of Algorithm 1 and some preliminary results for full pose (position in $\mathbb{R}^{3}$ and orientation in $\mathcal{SO}(3)$) motion planning in https://sites.google.com/view/lfd-neural-ode/home and the video submission.
@@ VII Conclusion & Future Work
APPENDIX
VII-A Full Pose Motions
We extend the CLF-NODE approach presented in Sections IV and V-B1 to learn full pose trajectories and generate stable motions in $\mathcal{SE}(3)$ space.
VII-A1 Neural ODE model
We extend the Neural ODE model (3) to learn both translational and rotational motions, i.e., the state $x(t)$ is the full-pose of the end effector. The state is represented as $x(t)=\begin{bmatrix}p(t),&q(t)\end{bmatrix}^{\top}$ where $p(t)\in\mathbb{R}^{3}$ is the Cartesian coordinates and $q(t)\in\mathcal{S}^{3}\subset\mathbb{R}^{4}$ lives on the 3-sphere $\mathcal{S}^{3}$ representing the rotation as a unit quaternion [43].
Quaternion Multiplication: We denote the operator $\otimes$ to represent the product between two quaternions, which is a linear operation [44] detailed as follows. Let $q=\begin{bmatrix}q_{1}&q_{2}&q_{3}&q_{4}\end{bmatrix}^{\top},r=\begin{bmatrix}r_{1}&r_{2}&r_{3}&r_{4}\end{bmatrix}^{\top}\in\mathbb{R}^{4}$ be two quaternions where the first elements $q_{1}$ and $r_{1}$ are the scalar parts, while the other three elements in each quaternion are the vector parts. Then, the left multiplication of $r$ by $q$ according to quaternion algebra [44] is $q\otimes r=L_{q}r$, where
$L_{q}=\begin{bmatrix}q_{1}&-q_{2}&-q_{3}&-q_{4}\\ q_{2}&q_{1}&-q_{4}&q_{3}\\ q_{3}&q_{4}&q_{1}&-q_{2}\\ q_{4}&-q_{3}&q_{2}&q_{1}\end{bmatrix}$ (15)
is the skew-symmetric matrix obtained from $q$. From (15), quaternion multiplication $\otimes$ is a linear operation, but non-commutative, i.e., $q\otimes r\neq r\otimes q$.
The state differential equation for full pose parameterized by the Neural ODE model is
$\frac{d\hat{x}(t)}{dt}=\begin{bmatrix}\frac{d\hat{p}(t)}{dt}\\ \frac{d\hat{q}(t)}{dt}\end{bmatrix}=\begin{bmatrix}f_{p}(\hat{x}(t))\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(\hat{x}(t))\end{bmatrix}\otimes\hat{q}(t)\right)\end{bmatrix}$, (16)
where hat $\hat{(\cdot)}$ explicitly represents the states predicted by the Neural ODE (16). The function $f_{p}:\mathbb{R}^{3}\times\mathcal{S}^{3}\to\mathbb{R}^{3}$ predicts the translational velocity $\frac{d\hat{p}(t)}{dt}$, while $f_{q}:\mathbb{R}^{3}\times\mathcal{S}^{3}\to\mathbb{R}^{3}$ predicts the angular velocity. Note that angular velocity is not the same as the rate of change of quaternions $\frac{d\hat{q}(t)}{dt}$. Given the demonstrations of full pose trajectories, we solve the empirical risk minimization problem (4) to learn the parameters $\theta$ of the neural network $f_{\theta}(\hat{x}(t))=\begin{bmatrix}f_{p}(\hat{x}(t)),&f_{q}(\hat{x}(t))\end{bmatrix}^{\top}$ where $f_{\theta}:\mathbb{R}^{3}\times\mathcal{S}^{3}\to\mathbb{R}^{6}$.
Remark: We use quaternions since they offer a compact, non-singular representation of orientation that avoids gimbal lock and allows for faster computation compared to angle-based representations or rotation matrices in $\mathcal{SO}(3)$. Moreover, the quaternion trajectory $\hat{q}(t)$ remains a unit quaternion throughout the integration process, provided the initial condition is a unit quaternion and the numerical integration is sufficiently accurate.
VII-A2 Virtual Control for Stable Motions
We build on Section V-B to generate stable full pose motions as
$\dot{x}=\begin{bmatrix}f_{p}(x)+u_{p}(x)\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(x)+u_{q}(x)\end{bmatrix}\otimes q\right)\end{bmatrix}$, (17)
where the additive virtual control term for linear velocity is $u_{p}(x)\in\mathbb{R}^{3}$, and that for angular velocity is $u_{q}(x)\in\mathbb{R}^{3}$. In our framework, we ensure convergence of the trajectory $x(t)$ generated by (17) to a target trajectory $x^{*}(t)=\begin{bmatrix}p^{*}(t),&q^{*}(t)\end{bmatrix}^{\top}\in\mathbb{R}^{3}\times\mathcal{S}^{3}$} by integrating Neural ODEs for complex motions and CLFs defined on the error between the current state and target. We begin by formulating the error dynamics as a control affine dynamical system (5), where the control inputs are the additive linear and angular velocity terms. Subsequently, we employ CLFs, as introduced in Section V-B1, to compute the virtual control input by solving a Quadratic Program (QP).
Error Dynamics: Given a target trajectory $x^{*}=\begin{bmatrix}p^{*},&q^{*}\end{bmatrix}^{\top}$, we define the error in position as $e_{p}=p-p^{*}$ and the error in quaternion as $e_{q}=q-q^{*}$, where we eliminate the explicit dependence on time $t$ for brevity. Although there are multiple error functions to quantify the distance between quaternions [45], the simple euclidean distance metric performed well in practice. The explicit control affine form of the full pose error dynamics is given below.
Proposition 1
Given a full pose target trajectory $x^{*}$, the error dynamics is a nonlinear control affine dynamical system
$\dot{e}=\begin{bmatrix}\dot{e}_{p}\\ \dot{e}_{q}\end{bmatrix}=\textbf{g}+\textbf{h}u(x),$ (18)
where the error is between the full pose current and target state:
$e=x-x^{*}=\begin{bmatrix}e_{p}\\ e_{q}\end{bmatrix}=\begin{bmatrix}p-p^{*}\\ q-q^{*}\end{bmatrix}$, (19)
the control input is $u(x)=\begin{bmatrix}u_{p}(x),&u_{q}(x)\end{bmatrix}^{\top}$, and
$\textbf{g}=\begin{bmatrix}f_{p}(x)-\dot{p}^{*}\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(x)\end{bmatrix}\otimes q\right)-\dot{q}^{*}\end{bmatrix}$, (20)
$\textbf{h}=\begin{bmatrix}\textbf{I}_{3}&\textbf{0}_{3\times 3}\\ \textbf{0}_{4\times 3}&0.5\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}L_{\overline{q}}\begin{bmatrix}\textbf{0}_{3}^{\top}\\ -\textbf{I}_{3}\end{bmatrix}\end{bmatrix}$, (21)
where $\overline{q}$ is the conjugate of the quaternion $q$ with $L_{\overline{q}}$ computed using (15), $\textbf{I}_{n}$ is an identity matrix of size $n$, $\textbf{0}_{m\times n}$ is a zero matrix of size $m\times n$, and $\textbf{0}_{n}$ is an n-vector of all zeros.
Proof:
From (17) and defining the error in position as $e_{p}=p-p^{*}$} and the error in quaternion as $e_{q}=q-q^{*}$, the error dynamics is given by
$\dot{e}=\begin{bmatrix}\dot{e}_{p}\\ \dot{e_{q}}\end{bmatrix}=\begin{bmatrix}f_{p}(x)+u_{p}(x)-\dot{p}^{*}\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(x)+u_{q}(x)\end{bmatrix}\otimes q\right)-\dot{q}^{*}\end{bmatrix}$. (22)
Since quaternion multiplication is linear (15), we can rewrite (22) as follows.
$\dot{e}=\begin{bmatrix}f_{p}(x)-\dot{p}^{*}\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(x)\end{bmatrix}\otimes q\right)-\dot{q}^{*}\end{bmatrix}+\begin{bmatrix}u_{p}(x)\\ 0.5\left(\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}\otimes q\right)\end{bmatrix}$. (23)
We introduce the quaternion $y=0.5\left(\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}\otimes q\right)$ and rewrite $y$ using quaternion conjugates.
Quaternion conjugate: Given a quaternion $q=\begin{bmatrix}q_{1}&q_{2}&q_{3}&q_{4}\end{bmatrix}^{\top}$, the complex conjugate of $q$ is
$\overline{q}=\begin{bmatrix}q_{1}&-q_{2}&-q_{3}&-q_{4}\end{bmatrix}^{\top}=\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}q.$ (24)
Since taking conjugates reverses the order of quaternion multiplication,
$y$ $=\overline{\overline{y}}=\overline{\overline{0.5\left(\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}\otimes q\right)}}=\overline{0.5\left(\overline{q}\otimes\overline{\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}}\right)}$.
Using (24),
$y$ $=\overline{0.5\left(\overline{q}\otimes\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}\right)}$,
and from (15),
$y$ $=\overline{0.5\left(L_{\overline{q}}\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}\right)}$,
and again using (24),
$y$ $=0.5\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}L_{\overline{q}}\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}\begin{bmatrix}0\\ u_{q}(x)\end{bmatrix}$
$\Rightarrow y$ $=0.5\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}L_{\overline{q}}\begin{bmatrix}\textbf{0}_{3}^{\top}\\ -\textbf{I}_{3}\end{bmatrix}u_{q}(x).$ (25)
From (25) and (23),
$\dot{e}$ $=\begin{bmatrix}f_{p}(x)-\dot{p}^{*}\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(x)\end{bmatrix}\otimes q\right)-\dot{q}^{*}\end{bmatrix}$+
$\begin{bmatrix}u_{p}(x)\\ 0.5\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}L_{\overline{q}}\begin{bmatrix}\textbf{0}_{3}^{\top}\\ -\textbf{I}_{3}\end{bmatrix}u_{q}(x)\end{bmatrix}$,
$\Rightarrow\dot{e}$ $=\begin{bmatrix}f_{p}(x)-\dot{p}^{*}\\ 0.5\left(\begin{bmatrix}0\\ f_{q}(x)\end{bmatrix}\otimes q\right)-\dot{q}^{*}\end{bmatrix}$+
$\begin{bmatrix}\textbf{I}_{3}&\textbf{0}_{3\times 3}\\ \textbf{0}_{4\times 3}&0.5\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}L_{\overline{q}}\begin{bmatrix}\textbf{0}_{3}^{\top}\\ -\textbf{I}_{3}\end{bmatrix}\end{bmatrix}\begin{bmatrix}u_{p}(x)\\ u_{q}(x)\end{bmatrix}$.
Based on the control-affine form of the error dynamics derived in Proposition 1, Theorem 1 implies that if there exists a Control Lyapunov Function (CLF) $V(\cdot)=-B(\cdot)$ for the full pose error dynamics (18), then, any feedback virtual control law $u(\cdot)$ that satisfies
$\nabla_{e}V(e)^{\top}\left(\textbf{g}+\textbf{h}u(x)\right)\leq-\alpha(V(e)),\ \forall\ e\in\mathbb{R}^{7}$ (26)
will drive the full pose error asymptotically to zero, where $\alpha(\cdot)$ is a class $\mathcal{K}_{\infty}$ function for CLF. During online motion planning, given the current full pose state of the robot $x$ and information about the full pose target trajectory $x^{*}$, we compute the minimal control effort $u(x)$ that satisfies (26) by setting
$u(x)=\operatorname*{arg\,min}_{v}\quad\bigl{\|}v\bigr{\|}^{2}_{2}$ (27)
[…]
$\left.\begin{bmatrix}\textbf{I}_{3}&\textbf{0}_{3\times 3}\\ \textbf{0}_{4\times 3}&0.5\begin{bmatrix}1&\textbf{0}_{3}^{\top}\\ \textbf{0}_{3}&-\textbf{I}_{3}\end{bmatrix}L_{\overline{q}}\begin{bmatrix}\textbf{0}_{3}^{\top}\\ -\textbf{I}_{3}\end{bmatrix}\end{bmatrix}v\right)\leq-\alpha(V(e)),$
where $\alpha(\cdot)$ defines how aggressively the robot tracks the target trajectory. The Lyapunov function we use is $V(e)=\|e\|_{2}^{2}$, but, note that any positive definite function is valid. Similar to (12), Problem (27) is also a Quadratic Program (QP), allowing for efficient computation in motion planning. This formulation stems from the control affine structure of the error dynamics derived in Proposition 1, which itself follows from the linearity of the quaternion multiplication operator (15). Solving (27) requires knowledge of the target point $x^{*}$, as detailed in the explanation below.
Choosing a Target Point: We use the same method described in Section V-B3 to choose the target position $p^{*}$ and quaternion $q^{*}$, where the target trajectory $x^{*}(t)$ is obtained by integrating the learned full pose Neural ODE model (16). In the closeness measure used to identify the nearest target point $m$ in Algorithm 1, we consider only the current position $p$, excluding the quaternion $q$. This is because, in our approach, the position and orientation are coupled through the full-pose Neural ODE model (16), such that the closest position target implicitly determines the appropriate quaternion target. This avoids the need for a separate proximity metric for quaternions, while leveraging the computational simplicity of the Euclidean distance in position space. Since the target point is obtained using the NODE model at each planning step, the target velocities are
$\dot{p}^{*}=f_{p}(x^{*}),\ \dot{q}^{*}=0.5\left(\begin{bmatrix}0\\ f_{q}(x^{*})\end{bmatrix}\otimes q^{*}\right),$ (28)
which are used in solving (27).
(a) DTW distance
(b) Quaternion error
(c) Training time
(d) Execution time
Figure 7: Comparison of performance metrics between CHN-sNODE and our approach (NODE) for reproducing full pose trajectories from the real robot dataset in [46].
VII-A3 Results
We validate our approach on the real robot dataset presented in [46] that contains 9 full pose motions representing different manipulation tasks. In Fig. 7, we compare our approach (NODE) with the CHN-sNODE method proposed in [46] across four metrics on the full-pose robotic tasks. NODE achieves comparable DTW distance and Quaternion error, indicating better trajectory alignment in both the position and rotational space. Moreover, NODE significantly reduces training time by over $50\%$ on average in the log scale, and improves execution time by several orders of magnitude, making it far more suitable for real-time applications.
(a) DoubleBendedLine
(b) PShape
(c) Snake
(d) MultiModels3
Figure 8: Trajectories predicted from the LASA test data set. The dashed black trajectories are from the data set. The solid green trajectories are predicted by the model. 1st row: Prediction from SEDS. 2nd row: Predictions from LPV-DS. 3rd row: Predictions by our Neural ODE (NODE) model. 4th row: DTW distance comparison on train data. 5th row: DTW distance comparison on test data.
VII-B Further Validation on Complex Trajectories
In Fig. 8, we compare the performance of our NODE model on four highly non-linear LASA trajectories where existing approaches: SEDS [3] and LPV-DS[16]; either directly go to the target point or do not capture the entire shape. We also present the trajectories predicted by our model on other non-linear and multi modal shapes in Fig. 9. We present obstacle avoidance scenarios in Fig. 11. We model one obstacle as a circle in Figs. 10(a) and 10(b), where, the region outside the blue obstacle is the safe region $\mathcal{C}$ as defined for the Control Barrier Function (CBF). The safe region is described using the CBF $B(x)=\|x-c\|_{2}^{2}-r^{2}$, where $c$ and $r$ is the center and radius of the obstacle, respectively. The other obstacle is concave and the barrier function of the concave obstacle is given by the union of two ellipses as below.
$B(x_{1},x_{2})=$ $\left(\frac{(x_{1}-x_{c1,1})^{2}}{a_{1}^{2}}+\frac{(x_{2}-x_{c1,2})^{2}}{b_{1}^{2}}-1\right)$ (29)
$\left(\frac{(x_{1}-x_{c2,1})^{2}}{a_{2}^{2}}+\frac{(x_{2}-x_{c2,2})^{2}}{b_{2}^{2}}-1\right),$
where $x=(x_{1},x_{2})$ is the Cartesian coordinates defined to be the state space. The parameters of the ellipses in Fig. 10(a) are $(x_{c1,1},x_{c1,2})=(-20,-10),a_{1}=5,b_{1}=7,(x_{c2,1},x_{c2,2})=(-23,-13),a_{2}=6,b_{2}=3$; and the parameters of the ellipses in Fig. 10(b) are $(x_{c1,1},x_{c1,2})=(-10,-22),a_{1}=3.5,b_{1}=5.5,(x_{c2,1},x_{c2,2})=(-8,-16),a_{2}=6,b_{2}=3$. The barrier function $B$ reshapes the vector field of the motion plan using the virtual control input $u(x)$ so that the robot motion always avoids the obstacle defined by the unsafe set: $\{x|B(x)<0\}.$
In Fig. 11, we present two scenarios where the robot switches from one mode to another during task execution. We use a linear form for the class $\mathcal{K}$ functions in (14): $\alpha(x)=k_{L}x,\gamma(x)=k_{B}x$; where, $k_{L}=4$} and $k_{B}=3$} are appropriate numerical constants for the CLF and CBF constraint, respectively. We note that we did not have to re-tune $k_{L}$ and $k_{B}$ for different LASA shapes. If we increase the value of $k_{L}$, then the trajectory tracking behavior is more aggressive. Similarly, if we increase the value of $k_{B}$, the the obstacle avoidance behaviour is less conservative. The trade off for more aggressive maneuvers is the increase in the magnitude of the commanded desired velocity – $f(x)+u(x)$ – which might lead to actuation limits in real robot experiments. We used $\lambda=0.5$ as the penalty term for the CLF-CBF QP. The forward looking horizon used is $N=3$ for Algorithm 1.
In Fig. 12, we compare periodic motions predicted by our method (NODE) with Imitation Flow (IFlow) [17] and Gaussian Process (GP) [11] based approach. In addition to the R shape presented in Fig. 5(d), we also present trajectory predictions for the I, O and S shapes from [17] in Figs. 12(a), 12(b) and 12(c). We compare the trajectory reproductions for 3D periodic wiping motions in Figs. 12(d) and 12(e), where we collect the training data by kinesthetically teaching the Franka robot the wiping tasks, similar to the spiral wiping motion in Fig. 5(e).
(a) BendedLine
(b) GShape
(c) heee
(d) NShape
(e) Leaf-1
(f) Leaf-2
(g) JShape-2
(h) Spoon
(i) CShape
(j) MultiModels1
(k) MultiModels2
(l) MultiModels4
Figure 9: Trajectories predicted by the NODE model without any safety or stability constraints on some non-linear and multi modal trajectories from the LASA handwriting data set. The dashed black trajectories are from the data set. The solid green trajectories are predicted by our model.
(a) Spoon with obstacles.
(b) Spoon with tilted obstacles
Figure 10: Stable and safe motion plan generated by our CBF-CLF NODE approach. Brown dot is the initial point, blue dot is the target point, and the light blue region are the obstacles.
(c) MultiModels2
(d) MultiModels3
Figure 11: Motion plan generated by our method when switching between modes. Brown dot is the initial point, and blue dot is the target point.
(a) I shape
(b) O shape
(c) S shape
(d) Wiping a whiteboard
(e) Wiping a mannequin
Figure 12: Comparison of trajectory reproductions between IFlow, GP and our approach (NODE) on more periodic trajectories.